# CookAR: Affordance Augmentations in Wearable AR to Support Kitchen Tool Interactions for People with Low Vision

### v1

@@ 3.1. Data Collection and Annotation
To train an effective AI model for affordance segmentation, we first need an appropriate dataset. However, to the best of our knowledge, there is no prior cooking tool dataset with annotations to enable affordance segmentation and recognition. Below, we describe our multi-step process to collect and annotate object affordances in egocentric cooking images. This labeled dataset is one contribution of our work and will be open-sourced upon paper acceptance to enable future research.
@@
Data Collection. To build a kitchen tool affordance image dataset, we used an egocentric video dataset called Epic Kitchens (Damen et al., 2018), which consists of 100 hours of video footage of sighted people cooking in their homes. We selected the Epic Kitchens video dataset since it not only involves a wide range of cooking scenarios with various kitchen tools, but also captures video feeds from the first-person perspective, which aligns with the egocentric nature of head-worn AR devices.
@@
Because the Epic dataset is so large, we needed to filter critical frames of interest. We used YOLOv8 (Jocher et al., 2023) trained on the MS COCO dataset (Lin et al., 2014) to detect and collect frames featuring cooking-related objects, such as spoons, knives, forks, cups, scissors, sinks, and dining tables. We skipped 20 frames after finding at least one of those objects to minimize repetition. We then manually reviewed the selected frames to empirically remove similar, excessively blurry, or irrelevant images, resulting in 4,928 key frames.
@@
Data Annotation & Augmentation. We then labeled these frames using the image annotation feature in Roboflow333https://roboflow.com, a platform that facilitates building and deploying end-to-end computer vision models by offering tools for annotation, training, and optimization. Roboflow enables labeling automation using the Segment Anything model (SAM) (Kirillov et al., 2023), allowing us to easily select and segment the interactive parts of objects (e.g., knife blade vs. knife handle) and add corresponding class labels.
@@
Drawing on prior work in low-vision cooking (Li et al., 2021, 2024; Wang et al., 2023b), the research team collectively decided on 18 distinct classes: Knife Blade, Knife Handle, Spoon Bowl, Spoon Handle, Fork Tines, Fork Handle, Scissor Blade, Scissor Handle, Ladle Bowl, Ladle Handle, Spatula Head, Spatula Handle, Pan Base, Pan Handle, Cup Base, Cup Handle, Carafe Base, and Carafe Handle (Figure 2). When labeling, we adhered to the following heuristic: (1) the object should visually resemble the class it is being labeled as; and (2) the object should serve functions similar to those of the label class. For instance, a large wooden spoon can be tricky to label, as it can resemble a spoon, ladle, or spatula, and have versatile use such as stirring a pan (like a spoon), scooping contents from a pot (like a ladle), or lifting eggs (like a spatula) across different images in the dataset. We labeled these ambiguous objects based on their shape and use in a given frame. The annotations were cross-checked between six members of the research team to reduce errors and bias.
@@
After annotating, we used various image augmentation techniques available on Roboflow to enhance the dataset for better generalizability across real-world scenarios, including: cropping with 0% minimum zoom and 40% maximum zoom, rotation between -15° and +15°, brightness between -15% and +15%, blur up to 2.5px, and noise up to 0.1% of pixels. We then adjusted the images to fit a 640x480 resolution (i.e., the MS COCO average image resolution (Lin et al., 2014)) to accommodate our chosen base model’s preferences and facilitate their use in future research. This resulted in a final dataset of 10,152 images. The entire dataset is publicly available here: https://github.com/makeabilitylab/CookAR.
@@ 3.2. Model Fine-Tuning
To provide real-time object affordance information to LV users, we fine-tuned the RTMDet model (Lyu et al., 2022), specifically its RTMDet-Ins-l variant, on our kitchen tool affordance dataset. This model is the current state-of-the-art in real-time instance segmentation444https://paperswithcode.com/sota/real-time-instance-segmentation-on-mscoco, offering robust accuracy and 300+ FPS on an NVIDIA 3090 GPU. RTMDet features large kernel depthwise convolution and batch normalization layers, pre-trained on MS COCO.
We opted to fine-tune RTMDet instead of training it from scratch, which allowed us to better leverage our smaller, class-specific dataset. To achieve this, we initialized the base RTMDet-Ins-l model with pre-trained weights, froze its backbone, and adjusted the model configuration file for our label classes, before training it on our kitchen tool affordance dataset. This customized model, dubbed RTMDet-Ins-l-Cook, was trained over 150 epochs with a batch size of 4 on a single CUDA-enabled NVIDIA 4080 GPU. Because RTMDet-Ins-l-Cook underwent fine-tuning on a dataset with affordance annotations, it can mimic an affordance segmentation model’s capabilities while retaining RTMDet’s real-time performance. We provide a technical evaluation in Section 4.
@@ 3.3. The CookAR Prototype
With our fine-tuned RTMDet-Ins-l-Cook model, we built CookAR, a wearable AR prototype that can recognize and visually augment the affordances of kitchen tools in real-time. To implement CookAR, we needed to address key technical and HCI challenges, including: (1) how to spatially highlight object affordances in 3D space; (2) how to run a real-time model capable of providing affordance feedback with minimal latency; (3) and how to best visually indicate affordances to LV users to support but not overwhelm their existing perceptual vision systems.
To generate visual augmentations that align with the object parts in a 3D space, we built a custom stereo video see-through AR system by combining the ZED Mini stereo camera with an Oculus Quest 2 VR headset, as off-the-shelf AR headsets (e.g., Microsoft HoloLens 2) do not yet support long range real-time depth sensing and heavy CV models. The affordance visualizations themselves are rendered as colored polygon overlays. Specifically, we used green (hexcode #3BE8B0) to indicate a graspable area and red (#FC626A) to indicate a risky area. See Figure 1. As our user study section describes (Section 5), we also further explored and brainstormed other affordance augmentation designs.
Because our real-time CV model is computationally expensive, CookAR is tethered to a laptop with a NVIDIA 4080 mobile GPU. The CookAR system first captures image frames using the ZED Mini stereo camera and streams them to an external server via the Transmission Control Protocol (TCP) for affordance segmentation by our RTMDet-Ins-l-Cook model (confidence threshold of 0.4). Then, it converts the resulting JSON with affordance masks and labels into a Protocol Buffers message555https://protobuf.dev for efficient streaming. This message is then sent back to be processed by the ZED Mini API (Stereolabs, 2024), which deserializes the message back into a JSON and creates a ZED-compatible texture (or colored overlay) for each affordance mask. Finally, the ZED Mini performs stereo depth estimation and overlays the textures onto the left and right image frames for binocular vision in the Quest 2 headset. While the ZED Mini stereo camera allowed us to prototype CookAR, it required our system to be tethered to a computer, which we alleviated by using lengthy cables.
In our latency analysis, we ran CookAR for five minutes and computed the average latency of each component: video streaming from ZED to computer took 16.76ms; affordance recognition took 15.95ms; result streaming back to ZED took 10.43ms; and depth estimation and augmentation rendering took 3.39ms. All other components had negligible impact on runtime. The overall latency is on average 46.82ms per frame (21.36 FPS), resulting in a near real-time system. System structure is summarized in Figure 3.
@@ 4. Technical Evaluation
We first conducted a technical evaluation of our fine-tuned RTMDet-Ins-l-Cook model, comparing its performance against the base RTMDet-Ins-l model on our kitchen tool affordance dataset. Findings indicate that our fine-tuned model is significantly more accurate in recognizing and segmenting affordances of cooking than the unmodified model.
@@ 4.1. Methods
To assess the performance of the base and fine-tuned models on our kitchen tool affordance dataset, we leveraged the model testing pipeline provided by the MMDetection library (Chen et al., 2019), a PyTorch-based open-source toolkit for object detection and segmentation, which performs evaluations using the test subset of a given dataset. Using Roboflow, we generated a testing set of 596 images with an 82-12-6 train-validation-test split and ensured that our fine-tuned model was not exposed to images in the test subset.
For instance segmentation tasks, accuracy is conventionally measured using three key metrics: segmentation mean average precision (mAP), AP at a 50% Intersection over Union (IOU) threshold (AP@50), and AP at a 75% IOU threshold (AP@75) (He et al., 2017). Each metric serves a distinct purpose:
mAP offers a comprehensive assessment of a model’s precision across various detection thresholds, by averaging precision at multiple recall levels for each class. It also aggregates results across a range of Intersection over Union (IoU) thresholds, from 0.5 to 0.95 in steps of 0.05, providing a holistic view of model performance across different degrees of overlap between the predicted masks and the ground truth;
AP@50 focuses on the precision of segmentation at a specific IoU threshold of 50%, a more lenient measure that considers a prediction correct if the overlap with the ground truth is at least half;
AP@75 evaluates precision at a stricter IoU threshold of 75%, demanding higher accuracy in the overlap between the predicted segmentation and the ground truth for a positive assessment.
IoU, central to these metrics, quantitatively evaluates the overlap between predicted segmentation masks and the actual ground truth, serving as a direct measure of accuracy in spatial alignment. We applied these metrics to compare the performance of our fine-tuned model against the baseline, aiming to capture the nuances of improvement across different levels of strictness in segmentation accuracy. We computed and compared the three metrics across the base and fine-tuned models.
@@ 4.2. Results
Our findings (Table 1) show that the base RTMDet-Ins-l model performs poorly on our affordance-specific dataset despite its competency on the COCO dataset (Lin et al., 2014) of 43.7% segmentation mAP. Conversely, our fine-tuned RTMDet-Ins-l-Cook model excelled in identifying and segmenting cooking tool components, demonstrating a significantly higher segmentation mAP of 46.3%, compared to the base model’s 12.3%. This improvement was also evident in our model’s performance at different IoU thresholds, with AP@50 and AP@75 reaching 74.9% and 48.6%, significantly outperforming the base model’s 19.9% and 13.2%. These results highlight the enhanced accuracy of our fine-tuned RTMDet-Ins-l-Cook model on our kitchen tool affordance dataset.
In addition, we show several inference results of RTMDet-Ins-l-Cook on images from the test subset of our dataset in Figure 4. Our model demonstrates impressive robustness, identifying and segmenting graspable, safe areas even when hands or other partial occlusions are present in the images.
@@ 5. User Study
We then evaluated our CookAR prototype in a three-part qualitative lab study with 10 low vision participants. As initial work, our goals were to evaluate how LV participants might benefit from a real-time object affordance augmentations to complete cooking tasks, solicit their reactions to a fully-functional but early-stage prototype (e.g., how do they react to augmentation errors), and co-brainstorm visual overlay designs via design probes. Participants provided feedback throughout the study and answered open-ended questions regarding their experiences, which were recorded and transcribed for later analysis.
5.1. Participants
To achieve a diverse participant pool, we recruited 10 LV participants from two different cities via mailing lists and snowball sampling. Participants were screened using a demographic questionnaire, which gathered information on age, gender, vision condition, and prior experience with AR and AI technologies. The average age was 62.2 years (), with a gender distribution of 70% female and 30% male. Participants had a broad range of low vision conditions with visual acuity ranging from 20/40 to 20/400 and visual field loss at different areas—see Table 2. Most participants reported little to no experience with AR and AI, except for P1 who had used both technologies.
@@ 5.1. Participants
| P9 | Female | 30 | Light Perception | 20/80 | Can make out faces with right eye. Left eye is blind. Visual field intact. |
@@ 5.2. Apparatus
The study was conducted in a well-lit lab environment. Participants sat in front of a large table, where we placed nine different kitchen tools—knife, spoon, fork, scissors, ladle, spatula, pan/pot, cup, and carafe. We used a dark green table cloth for the experiment table to produce low contrast. We also prepared a yellow wooden cutting board, a bowl, a piece of cheese, and a stick of butter for the participants to use in the study, however, CookAR can only recognize and augment the nine aforementioned kitchen tools at the current stage. Lastly, we recorded the experiment using a separate laptop and a smartphone on a tripod.
@@ 5.3. Procedure
The single-session 90-minute study consisted of three phases. In Part 1, we asked participants to grab cooking utensils in three conditions: (1) without CookAR, (2) with CookAR displaying augmentations for whole objects, and (3) with CookAR displaying affordance augmentations (Figure 5). In Part 2, participants completed a full cooking task where they made macaroni and cheese while using CookAR. Finally, in Part 3, participants brainstormed tool affordances and desired augmentation designs while examining and critiquing design probes. Prior to the study tasks, participants completed a tutorial to become more familiar with CookAR. They then interacted with our AR device for 30 minutes: a 5-minute tutorial, 10 minutes of object grabbing, and 15 minutes of free-form cooking. We provide more details below.
@@
Tutorial. Participants were asked to first complete a tutorial task: interact with a cooking pot using CookAR. Participants wore and adjusted the Oculus Quest 2 headset and became acquainted with CookAR and its affordance augmentations. Once participants achieved a comfortable fit and an understanding of CookAR, the study proceeded.
@@
Part 1: Comparison Study. In Part 1, participants were asked to locate and pick up cooking tools under three conditions: (1) without CookAR, (2) with CookAR displaying whole object augmentations, and (3) with CookAR displaying affordance augmentations (See Figure 5). We counterbalanced the condition order via Latin Square. In each condition, participants conducted five trials of tool-picking-up tasks. We randomly chose a cooking tool per trial from the nine kitchen tools on the experiment table (See Section 5.2). Participants were asked to keep their eyes closed until the researcher named an object to reduce the effect of memory on task performance. Moreover, the researcher rearranged the placement and angle of the cooking tools between each condition. After each condition, we asked participants three 7-point Likert questions about effectiveness, comfort, and distraction, as well as open-ended questions regarding their experience with each augmentation condition. After all 15 trials, we asked participants to compare the pros and cons of the three conditions, suggest improvements for CookAR, and identify potential applications of CookAR outside of kitchen contexts.
@@
Part 2: Full Cooking Task. In Part 2 of the study, we asked participants to cook a macaroni and cheese dish using ARCook with affordance augmentations. We provided participants with step-by-step instructions to ensure that participants interact with all nine objects ARCook can recognize: (1) grab a cup of water and pour it into a carafe, (2) pour water into a pot using a carafe, (3) cut a piece of butter using a knife, (4) cut a piece of cheese using a pair of scissors, (5) put the macaroni, butter, and cheese into a pot, (6) stir with a spoon, (7) stir with a spatula, (8) place the finished macaroni and cheese in a bowl using a ladle, and (9) pick up a fork and enjoy. For the safety of our participants, we supplied a knife with a dull edge and avoided the use of heat. As participants completed this task, they were encouraged to think aloud, articulating how the affordance augmentations support or hinder their activities, how CookAR impacted their overall cooking experience, and any suggestions they had for augmentation designs. After completing the free-form cooking task, we asked participants to reflect on these same topics through open-ended questions.
@@
Part 3: Brainstorming and Co-Design. In Part 3 of the study, we asked participants to brainstorm future designs and applications of CookAR. We first presented design probes (See Figure 6) and asked participants to critique them. The design probes illustrate various designs, including outlines which reduce visual clutter in comparison to solid-colored overlays, solely displaying either the grabbable or the hazardous augmentation, highlighting the more specific hazardous part such as the sharp edge of a knife blade rather than the whole blade, employing arrows to widen the area covered by the augmentations, and introducing a visual warning system when the user’s hand gets too close to a risky area. The design probes are inspired by prior work on low vision augmentations (Zhao et al., 2016, 2019; Fox et al., 2023; Lang and Machulla, 2021). Subsequently, we invited participants to propose augmentation design ideas for both simple objects like knifes and more intricate objects with more interactive parts beyond grabbable and hazardous areas, like a carafe with many holes and buttons. Lastly, we asked participants to identify scenarios other than a kitchen where CookAR might be beneficial.

### v2

@@ 3.1. Data Collection and Annotation
To train a CV model for affordance segmentation, we first needed a labeled dataset. However, to our knowledge, there is no prior cooking tool dataset with annotations to enable affordance segmentation. Below, we describe our multi-step process to collect and annotate object affordances in egocentric cooking images.
@@
Data Collection. To build our kitchen tool affordance image dataset, we used an egocentric video repository called Epic Kitchens (Damen et al., 2018), which consists of 100 hours of video footage of sighted people cooking in their homes. We selected this dataset since it not only involves a wide range of cooking scenarios with various kitchen tools but also captures video feeds from a first-person perspective, which aligns with the egocentric nature of head-worn AR devices.
@@
Because the Epic Kitchens dataset is large, we first needed to filter for frames of interest. We used YOLOv8 (Jocher et al., 2023) trained on the MS COCO dataset (Lin et al., 2014) to detect and collect frames featuring cooking-related objects, such as spoons, knives, forks, cups, scissors, sinks, and dining tables. To minimize repetition, we skipped 20 frames after finding at least one of those objects. We then manually reviewed the selected frames to empirically remove similar, excessively blurry, or irrelevant images, resulting in 4,928 key frames.
@@
Data Annotation & Augmentation. We then labeled these frames using Roboflow333https://roboflow.com, an online tool for annotating, training, and optimizing CV models. Roboflow also supports labeling automation using the Segment Anything model (SAM) (Kirillov et al., 2023), allowing us to easily select and segment interactive parts of objects (e.g., knife blade vs. knife handle) and add corresponding class labels.
@@
Drawing on prior research (Li et al., 2021, 2024; Wang et al., 2023b), the research team identified 18 distinct classes of kitchen tools commonly used by BLV individuals: Knife Blade, Knife Handle, Spoon Bowl, Spoon Handle, Fork Tines, Fork Handle, Scissor Blade, Scissor Handle, Ladle Bowl, Ladle Handle, Spatula Head, Spatula Handle, Pan Base, Pan Handle, Cup Base, Cup Handle, Carafe Base, and Carafe Handle (Figure 2). When labeling, we adhered to the following heuristic: (1) the object should visually resemble the class it is labeled as; and (2) the object should serve functions similar to those of the label class. For instance, a large wooden spoon can be tricky to label, as it can resemble a spoon, ladle, or spatula, and have versatile use such as stirring a pan (like a spoon), scooping contents from a pot (like a ladle), or lifting eggs (like a spatula) across different images in the dataset. We labeled these ambiguous objects based on their shape and use in a given frame. Six research team members performed the annotations, each labeling a subset of images and having their work reviewed by another researcher to reduce errors and bias.
@@
After annotating, we used various image augmentation techniques available on Roboflow to enhance the dataset for better generalizability across real-world scenarios, including: cropping with 0% minimum zoom and 40% maximum zoom, rotation between -15° and +15°, brightness between -15% and +15%, blur up to 2.5px, and noise up to 0.1% of pixels. We then adjusted the images to fit a 640x480 resolution (i.e., the MS COCO average image resolution (Lin et al., 2014)) to accommodate our chosen base model’s preferences and facilitate their use in future research. This resulted in a final dataset of 10,152 images.
@@ 3.2. Model Fine-Tuning
To provide real-time object affordance information to LV users, we fine-tuned the RTMDet model (Lyu et al., 2022), specifically its RTMDet-Ins-l variant, on our kitchen tool affordance dataset. This model is the current state-of-the-art in real-time instance segmentation444https://paperswithcode.com/sota/real-time-instance-segmentation-on-mscoco, offering robust accuracy and 300+ FPS on an NVIDIA 3090 GPU. RTMDet features large kernel depthwise convolution and batch normalization layers, pre-trained on MS COCO (Lin et al., 2014).
We opted to fine-tune RTMDet instead of training it from scratch, as this allowed us to better leverage our smaller, class-specific dataset. To achieve this, we leveraged the fine-tuning pipeline provided by the MMDetection library (Chen et al., 2019), a PyTorch-based open-source toolkit for object detection and segmentation, which supports various models including RTMDet. We initialized the base RTMDet-Ins-l model with pre-trained weights, froze its backbone, and adjusted the model configuration file for our label classes before training it on our kitchen tool affordance dataset. This customized model, dubbed RTMDet-Ins-l-Cook, was trained over 150 epochs with a batch size of 4 on a single CUDA-enabled NVIDIA 4080 GPU. Because RTMDet-Ins-l-Cook underwent fine-tuning on a dataset with affordance annotations, it can mimic an affordance segmentation model’s capabilities while retaining RTMDet’s real-time performance. We provide a technical evaluation of our model in Section 4. We also open-sourced our fine-tuning steps and code, giving researchers the tools to expand our dataset and fine-tune their own models: https://github.com/makeabilitylab/mmdet-fine-tuning.
@@ 3.3. The CookAR Prototype
With our RTMDet-Ins-l-Cook model, we built CookAR, a wearable AR prototype that can recognize and visually augment the affordances of kitchen tools in near real-time. To implement CookAR, we addressed three key technical and HCI challenges, including: (1) how to spatially highlight object affordances in 3D space; (2) how to develop a real-time system pipeline to provide affordance feedback with minimal latency; and (3) how to best visually indicate affordances to LV users to support but not overwhelm their existing visual perceptions. See video figure for a demonstration.
To generate visual augmentations that align with the object parts in 3D space (e.g., a knife grip, a cup handle), we built a custom stereo video see-through AR system by combining the ZED Mini stereo camera with an Oculus Quest 2 VR headset. While off-the-shelf AR headsets such as the Microsoft HoloLens 2 may eliminate the need for a bulky video passthrough system, they do not yet support long range real-time depth sensing. We visualize the affordance representations as colored polygon overlays. Though the colors are settable, we currently use green (hexcode #3BE8B0) to indicate a graspable area and red (#FC626A) to indicate a risky area (Figure 1). As described in our user study section, we also further explored and brainstormed other affordances and augmentation designs.
Because our real-time CV model is computationally expensive, CookAR is tethered to a laptop with a NVIDIA 4080 mobile GPU. The CookAR system first captures image frames using the ZED Mini stereo camera and streams them to an external server via the Transmission Control Protocol (TCP) for affordance segmentation by our RTMDet-Ins-l-Cook model (confidence threshold of 0.4). Then, the server converts the resulting JSON with affordance masks and labels into a Protocol Buffers message555https://protobuf.dev for efficient streaming. This message is then sent back for processing by the ZED Mini API (Stereolabs, 2024), which deserializes the message back into a JSON and creates a ZED-compatible texture (or colored overlay) for each affordance mask. Finally, the ZED Mini performs stereo depth estimation and overlays the textures onto the left and right image frames for binocular vision in the Quest 2 headset. To enable participants to move freely during the study session, we connected the CookAR system to a computer using long (16 feet) cables.
In our latency analysis, we ran CookAR for five minutes and computed the average latency of each component: video streaming from ZED to the server took 16.76ms; affordance recognition took 15.95ms; result streaming back to ZED took 10.43ms; and depth estimation and augmentation rendering took 3.39ms. All other components had negligible impact on runtime. The overall latency is on average 46.82ms per frame (21.36 FPS), resulting in a near real-time system. See Figure 3 for a system flow diagram.
@@ 4. Technical Evaluation
We first conducted a technical evaluation of our RTMDet-Ins-l-Cook model, comparing its performance against the base RTMDet-Ins-l model on our kitchen tool affordance dataset. Findings indicate that our model is significantly more accurate in recognizing and segmenting affordances of cooking tools than the unmodified model.
@@ 4.1. Methods
To assess the performance of the base and fine-tuned models on our kitchen tool affordance dataset, we used MMDetection’s (Chen et al., 2019) model testing pipeline, which performs evaluations using the test subset of a given dataset. With Roboflow, we generated a test set of 596 images with an 82-12-6 train-validation-test split and ensured that our model was not exposed to images in the test subset.
For instance segmentation tasks, accuracy is conventionally measured using three key metrics: segmentation mean average precision (mAP), AP at a 50% Intersection over Union (IOU) threshold (AP@50), and AP at a 75% IOU threshold (AP@75) (He et al., 2017). IoU, central to these metrics, quantitatively evaluates the overlap between predicted segmentation masks and the actual ground truth, serving as a direct measure of accuracy in spatial alignment. We explain each metric in detail below:
mAP offers a comprehensive assessment of a model’s precision across various detection thresholds by averaging precision at multiple recall levels for each class. It also aggregates results across a range of Intersection over Union (IoU) thresholds, from 0.5 to 0.95 in steps of 0.05, providing a holistic view of model performance across different degrees of overlap between the predicted masks and the ground truth;
AP@50 uses a precision of segmentation at the IoU threshold of 50%, a more lenient measure that considers a prediction correct if the overlap with ground truth is at least half;
AP@75 evaluates precision at a stricter IoU threshold of 75%, demanding higher accuracy in the overlap between the predicted segmentation and ground truth.
We applied these metrics to compare our model’s performance against the baseline, aiming to capture the nuances of improvement across different levels of strictness in segmentation accuracy.
@@ 4.2. Results
Our findings (Table 1) indicate that the base RTMDet-Ins-l model performs poorly on our affordance dataset despite its competency on the COCO dataset (Lin et al., 2014) with a 43.7% segmentation mAP. Moreover, our fine-tuned RTMDet-Ins-l-Cook model excelled in identifying and segmenting cooking tool components, demonstrating a significantly higher segmentation mAP of 46.3%, compared to the base model’s 12.3%. This improvement was also evident in our model’s performance at different IoU thresholds, with AP@50 and AP@75 reaching 74.9% and 48.6%, significantly outperforming the base model’s 19.9% and 13.2%. In Figure 4, we show several example inference results of RTMDet-Ins-l-Cook on test images. Our model demonstrates impressive robustness, identifying and segmenting graspable, safe areas even when hands or other partial occlusions are present in the images. Overall, these findings highlight the enhanced accuracy of our RTMDet-Ins-l-Cook model.
@@ 5. User Study
As a complement to our technical evaluation, we conducted a three-part qualitative lab study with 10 low vision (LV) participants. Our goals were threefold: first, to evaluate how LV participants might benefit from real-time object affordance augmentations when completing cooking tasks; second, to solicit their reactions to a fully-functional but early-stage prototype (e.g., how do they react to augmentation errors); finally, to co-brainstorm visual overlay designs via design probes. Participants provided feedback throughout the study and answered open-ended questions regarding their experiences, which were recorded and transcribed for later analysis.
@@
| P9 | Female | 30 | Light Perception | 20/80 | Can make out faces w/right eye. Left eye blind. Visual field intact. |
@@
5.1. Participants
To achieve a diverse participant pool, we recruited 10 LV participants from two cities (Madison, WI and Seattle, WA) via mailing lists and snowball sampling. Participants were screened using a demographic questionnaire, which gathered information on age, gender, vision condition, and prior experience with AR and AI technologies. The average age was 62.2 years (SD=19.6), with a gender distribution of 70% female and 30% male. Participants had a broad range of low vision conditions with visual acuity ranging from 20/40 to 20/400 and visual field loss at different areas—see Table 2. Most participants reported little to no experience with AR and AI, except for P1 who had used both technologies.
@@ 5.2. Apparatus
The study was conducted in a well-lit lab environment. Participants sat in front of a large table, where we placed nine different kitchen tools—knife, spoon, fork, scissors, ladle, spatula, pan/pot, cup, and carafe. We used a dark green table cloth to simulate a visually challenging environment with low contrast. We also prepared a yellow wooden cutting board, a bowl, a piece of cheese, and a stick of butter for the participants to use in the study, although CookAR can only recognize and augment the nine aforementioned kitchen tools at the current stage. Lastly, we recorded the experiment using a laptop and a smartphone on a tripod.
@@ 5.3. Procedure
The single-session 90-minute study consisted of three phases. In Part 1, we asked participants to grab cooking utensils with CookAR and two baselines. In Part 2, participants completed a full cooking task where they made macaroni and cheese while using CookAR with affordance augmentations. Finally, in Part 3, participants brainstormed tool affordances and desired augmentation designs while examining and critiquing design probes. Prior to the study tasks, participants completed a tutorial to become familiar with CookAR. In total, participants interacted with our AR device for 30 minutes across the study: a 5-minute tutorial, 10 minutes of object grabbing, and 15 minutes of free-form cooking. We provide more details below. The full study protocol is in the Appendix.
@@
Tutorial. Participants first completed a tutorial task, where they interacted with a cooking pot using CookAR. Participants wore and adjusted the Oculus Quest 2 headset and freely explored CookAR and its affordance augmentations. Once they achieved a comfortable fit and an understanding of CookAR, the study proceeded.
@@
Part 1: Tool Grabbing Task. In Part 1, participants were asked to locate and pick up cooking tools under three conditions: (1) real-world baseline (i.e., without CookAR), (2) augmentation baseline (i.e., whole object augmentations) (Zhao et al., 2016; Fox et al., 2023), and (3) CookAR (i.e., affordance augmentations) (See Figure 5). We counterbalanced the condition order via Latin Square. In each condition, participants conducted five trials of picking-up tool tasks. We randomly chose a cooking tool per trial from the nine kitchen tools on the experiment table (See Section 5.2). Participants were asked to keep their eyes closed until the researcher named an object to reduce the effect of memory on task performance. Moreover, the researcher rearranged the placement and angle of the cooking tools between each condition. After each condition, we asked participants three 7-point Likert questions about effectiveness, comfort, and distraction, as well as open-ended questions regarding their experience with each augmentation condition. After all 15 trials, we asked participants to compare the pros and cons of the three conditions, suggest improvements for CookAR, and identify potential applications of CookAR outside of kitchen contexts.
@@
Part 2: Full Cooking Task. In Part 2, we asked participants to cook a macaroni and cheese dish using CookAR with affordance augmentations. We provided participants with step-by-step instructions for consistency and to ensure participants interacted with all nine objects CookAR can recognize: (1) grab a cup of water and pour it into a carafe, (2) pour water into a pot using a carafe, (3) cut a piece of butter using a knife, (4) cut a piece of cheese using a pair of scissors, (5) put the macaroni, butter, and cheese into a pot, (6) stir with a spoon, (7) stir with a spatula, (8) place the finished macaroni and cheese in a bowl using a ladle, and (9) pick up a fork and enjoy. For the safety of our participants, we supplied a knife with a dull edge and avoided the use of heat. As participants completed this task, they were encouraged to think aloud, articulating how the affordance augmentations supported or hindered their activities, how CookAR impacted their overall cooking experience, and any suggestions they had for augmentation designs. After completing the free-form cooking task, we asked participants to reflect on these same topics through open-ended questions.
@@
Part 3: Brainstorming and Co-Design. In Part 3, we asked participants to brainstorm future designs and applications of CookAR. Drawing on prior work in low vision augmentation (Zhao et al., 2016, 2019; Fox et al., 2023; Lang and Machulla, 2021), we created and presented design probes of various augmentation designs (Figure 6) and asked participants for feedback. The design probes included: (1) outlines to reduce visual clutter in comparison to solid-colored overlays, (2) solely displaying either the grabbable or the hazardous augmentation, (3) highlighting the more specific hazardous part such as the sharp edge of a knife blade rather than the whole blade, (4) employing arrows to widen the area covered by the augmentations, and (5) introducing a visual warning system when the user’s hand gets too close to a risky area. After presenting our own designs, we invited participants to propose other ideas for both simple objects like knives as well as more complex objects with additional interactive parts beyond grabbable and hazardous areas, like a carafe, which has many openings and buttons with different purposes. Lastly, we asked participants to identify other scenarios where affordance augmentations might be beneficial.